OpenStudy (anonymous):

(9a^6 b^6) over (6a^6 b^10) divided by (3a^4 b^12) over (a^2 b^4)

jhonyy9 (jhonyy9):

\[3/b ^{4}/3a ^{2} b ^{8}=3/b ^{4}*1/3a ^{2}b ^{8}= 1/a ^{2} b ^{12}\]